OpenStudy (anonymous):

Determine whether the geometric series is convergent or divergent. And if it is convergent, what is the sum?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (-3)^{n-1} }{ 4^{n} }\]

OpenStudy (anonymous):

i just don't know which value a and r is.

OpenStudy (amistre64):

what are our criteria for convergence? and the r will be the part that his the exponent

OpenStudy (amistre64):

\[\frac{(-3)^{n-1}}{4^n}\] \[\frac{(-1)^{n-1}~{3^n}~3^{-1}}{4^n}\] etc ...

OpenStudy (anonymous):

you kinda of lost me here.

OpenStudy (anonymous):

i know for convergence, r needs to be less than 1

OpenStudy (anonymous):

convergence is when r is between -1 and 1

OpenStudy (anonymous):

yeah, but im not sure how to figure ou how to find r.

OpenStudy (anonymous):

well let c n1= 1/4 right?

OpenStudy (anonymous):

?

OpenStudy (anonymous):